Average: /2 4. Problems and Applications 04 2023-21 Nimbus, Inc., makes brooms and then sells them door-to-door. Here is the relationship between the number of workers and Nimbuss output in a se Inform given day: Page Herel In the following table, complete the marginal product column. nfo Workers Output Marginal Product Total CostMarginal Cost Avge Total Cost e Quest 70 120 4 165 5 200 6 220 7 230 ig ed Ass and Re A worker costs $120 a day, and the firm has fixed costs of $180. Use this information to fill in the column for total cost in the previous table. inking Messa Fill n the column for marginal cost in the previous table. (xtint: M des Fill in the column for average total cost in the previous table. (Hint: ATC )

media%2F03a%2F03af7152-0add-4cf9-9068-a3
Answer #1
Workers Output Marginal product Variable cost Fixed cost Total cost Marginal cost Average total cost
0 0 - 120(0)=0 180 (0+180)=180 - -
1 30 30 120(1)=120 180 (120+180)=300 (300-180)/(30) = 4 (300/30)=10
2 70 40 (120)(2)=240 180 (240+180)=420 (420-300)/(70-30) =3 (420/70)=6
3 120 50 (120)(3)=360 180 (360+180)=540 (540-420)/(120-70) = 2.4 (540/120)=4.5
4 165 45 (120)(4)=480 180 (480+180)=660 (660-540)/(165-120)= 2.67 (660/165)= 4
5 200 35 (120)(5)=600 180 (600+180)=780 (780-660)/(200-165)= 3.43 (780/200)=3.9
6 220 20 (120)(6)=720 180 (720+180)=900 (900-780)/(220-200)= 6 (900/220)=4.09
7 230 10 (120)(7)=840 180 (840+180)=1020 (1020-900)/(230-220)= 12 (1020/230)=4.43

When marginal product is rising , marginal cost is falling. (As we can see from the above table).

When marginal cost is greater than average total cost , average total cost is rising. (As we can see from the above table).

Average total cost is U-shaped . (Earlier ATC is falling and then rising as pattern of ATC in table).

Hence, statement (A),(C), (D) are correct.

Statement (B) is not true because marginal cost is falling first and then rising.